OpenStudy (anonymous):

. Jeremy is building a large deck for a community center. The deck is shaped as a rectangle. The width of the deck is 29 feet. The perimeter of the deck is to be at least 134 feet. a. Write an inequality that represents all possible values for the length of the deck. b. Find all possible values for the length of the deck.

OpenStudy (texaschic101):

A. P = 2(W + L) 134 => 2(29 + L) 134 = > 58 + 2L 134 - 58 = > 2L 76 => 2L 76/2 = > L 38 = > L The length has to be equal to or greater then 38 ft. You might want to get a second opinion on this :/

OpenStudy (texaschic101):

wait...I messed up....The length has to be less then or equal to 38 ft.